### v1

# Unleashing the Potential of Diffusion Models for Incomplete Data Imputation

Abstract
This paper introduces DiffPuter, an iterative method for missing data imputation that leverages the Expectation-Maximization (EM) algorithm and Diffusion Models. By treating missing data as hidden variables that can be updated during model training, we frame the missing data imputation task as an EM problem. During the M-step, DiffPuter employs a diffusion model to learn the joint distribution of both the observed and currently estimated missing data. In the E-step, DiffPuter re-estimates the missing data based on the conditional probability given the observed data, utilizing the diffusion model learned in the M-step. Starting with an initial imputation, DiffPuter alternates between the M-step and E-step until convergence. Through this iterative process, DiffPuter progressively refines the complete data distribution, yielding increasingly accurate estimations of the missing data. Our theoretical analysis demonstrates that the unconditional training and conditional sampling processes of the diffusion model align precisely with the objectives of the M-step and E-step, respectively. Empirical evaluations across 10 diverse datasets and comparisons with 16 different imputation methods highlight DiffPuter’s superior performance. Notably, DiffPuter achieves an average improvement of 8.10% in MAE and 5.64% in RMSE compared to the most competitive existing method. The code is available at https://github.com/hengruizhang98/DiffPuter.
1 Introduction
In the field of data science and machine learning, missing data in tabular datasets is a common issue that can severely impair the performance of predictive models and the reliability of statistical analyses. Missing data can result from various factors, including data entry errors, non-responses in surveys, and system errors during data collection [1, 15, 4]. Properly handling missing data is essential, as improper treatment can lead to biased estimates, reduced statistical power, and invalid conclusions.
A plethora of work proposed over the past decades has propelled the development of missing data imputation research. Early classical methods often relied on partially observed statistical features to impute missing values or were based on shallow machine learning techniques, such as KNN [23], or simple parametric models, such as Bayesian models [17] or Gaussian Mixture Models [5]. These methods offer ample interpretability; however, they are constrained by capacity limitations, making it challenging to achieve satisfying performance. With the advent of deep learning, recent research has primarily focused on predictive [28, 29, 14] or generative deep models [30, 18, 25] for missing data imputation. Predictive models learn to predict the target entries conditioned on other observed entries, guided by masking mechanisms [3] or graph regularization techniques [31, 35]. By contrast, generative methods learn the joint distribution of missing entries and observed entries and try to impute the missing data via conditional sampling [25, 20, 18, 30, 21, 34]. Despite employing state-of-the-art generative models [12, 24, 6], generative-imputation methods still fall short compared to predictive methods. We conjecture that this is due to the following reasons: 1) Incomplete likelihood: generative models need to estimate the joint distribution of missing data and observed data. However, since the missing data itself is unknown, there is an inherent error in the estimated data density. 2) Conditional inference: even if we have a generative model that can generate samples from the complete data distribution faithfully, it is still challenging to perform conditional inference since the learned complete distribution either has very complex parametrization or is learned implicitly, and is hard to perform conditional inference, e.g., conditional sampling.
This paper introduces DiffPuter, a principled generative method for missing data imputation based on the Expectation-Maximization (EM) algorithm [2] and Diffusion Models [27, 10], designed to address the aforementioned issues. The EM algorithm is a well-established method for missing data imputation, capable of addressing the incomplete likelihood issue by iteratively refining the values of the missing data. It operates under the intuitive assumption that knowing the values of the missing variables simplifies the maximum likelihood problem for the imputation task. However, integrating the EM algorithm with deep generative models has been less explored, primarily due to the challenges of performing conditional inference for deep generative models without additional assumptions about distribution parameterization [20, 18, 30]. In this paper, we demonstrate that combining the diffusion model with the EM algorithm creates an effective imputation method for incomplete tabular data. Specifically: 1) In the M-step, DiffPuter employs a diffusion model to learn the joint distribution of the missing and observed data. 2) In the E-step, DiffPuter uses the learned diffusion model to perform flexible and accurate conditional sampling by mixing the forward process for observed entries with the reverse process for missing entries. Theoretically, we show that DiffPuter’s M-step corresponds to the maximum likelihood estimation of the data density, while its E-step represents the Expected A Posteriori (EAP) estimation of the missing values, conditioned on the observed values.
We conducted experiments on 10 benchmark tabular datasets containing both continuous and discrete features under various missing data scenarios. We compared the performance of DiffPuter with 16 competitive imputation methods from different categories. Experimental results demonstrate the superior performance of DiffPuter across all settings and on almost all datasets. In addition, experiments demonstrate that DiffPuter’s iterative training can effectively and gradually reduce the error in density estimation, thereby improving the performance of imputation. Ablation studies also demonstrate DiffPuter’s robustness to different missing mechanisms and missing ratios.
2 Related Works
Iterative Methods for Missing Data Imputation.
Iterative imputation is a widely used approach due to its ability to continuously refine predictions of missing data, resulting in more accurate imputation outcomes. This iterative process is especially crucial for methods requiring an initial estimation of the missing data. The Expectation-Maximization (EM) algorithm [2], a classical method, can be employed for missing data imputation. However, earlier applications often assume simple data distributions, such as mixtures of Gaussians for continuous data or Bernoulli and multinomial densities for discrete data [5]. These assumptions limit the imputation capabilities of these methods due to the restricted density estimation of simple distributions. The integration of EM with deep generative models remains underexplored. A closely related approach is MCFlow [25], which iteratively imputes missing data using normalizing flows [24]. However, MCFlow focuses on recovering missing data through maximum likelihood rather than expectation, and its conditional imputation is achieved through soft regularization instead of precise sampling based on the conditional distribution. Beyond EM, the concept of iterative training is prevalent in state-of-the-art deep learning-based imputation methods. For instance, IGRM [35] constructs a graph from all dataset samples and introduces the concept of friend networks, which are iteratively updated during the imputation learning process. HyperImpute [8] proposes an AutoML imputation method that iteratively refines both model selection and imputed values.
Diffusion Models for Missing Data Imputation.
We are not the first to utilize diffusion models for missing data imputation. TabCSDI [34] employs a conditional diffusion model to learn the distribution of masked observed entries conditioned on the unmasked observed entries. MissDiff [21] uses a diffusion model to learn the density of tabular data with missing values by masking the observed entries. Although MissDiff was not originally intended for imputation, it can be easily adapted for this task. Other methods, despite claiming applicability for imputation, are trained on complete data and evaluated on incomplete testing data [32, 9]. These approaches contradict the focus of this study, where the training data itself contains missing values. Additionally, all the aforementioned methods use one-step imputations, which overlook the issue that missing data in the training set can lead to inaccurate data density estimation.
3 Preliminaries
3.1 Missing Value Imputation for Incomplete Data
This paper addresses the missing value imputation task for incomplete data, where only partial data entries are observable during the training process. Formally, let the complete $d$-dimensional data be denoted as $\mathbf{x}\sim p_{\rm data}(\mathbf{x})\in{\mathbb{R}}^{d}$. For each data sample, $\mathbf{x}$, there is a binary mask $\mathbf{m}\in\{0,1\}^{d}$ indicating the location of missing entries for $\mathbf{x}$. Let the subscript $k$ denote the $k$-th entry of a vector, then $m_{k}=1$ stands for missing entries while $m_{k}=0$ stands for observable entries.
We further use $\mathbf{x}^{\rm obs}$ and $\mathbf{x}^{\rm mis}$ to denote the observed data and missing data, respectively (i.e., $\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$). Note that $\mathbf{x}^{\rm obs}$ is the fixed ground-truth observation, while $\mathbf{x}^{\rm miss}$ is conceptual, unknown, and we we aim to estimate it. The missing value imputation task aims to predict the missing entries $\mathbf{x}^{\rm mis}$ based on the observed entries $\mathbf{x}^{\rm obs}$.
In-sample vs. Out-of-sample imputation.
The missing data imputation task can be categorized into two types: in-sample and out-of-sample. In-sample imputation means that the model only has to impute the missing entries in the training set, while out-of-sample imputation requires the model’s capacity to generalize to the unseen data records without fitting its parameters again. Not all imputation methods can generalize to out-of-sample imputation tasks. For example, methods that directly treat the missing values as learnable parameters [19, 33] are hard to apply to unseen records. A desirable imputation method is expected to perform well on both in-sample and out-of-sample imputation tasks, and this paper studies both of the two settings.
3.2 Formulating Missing Data Imputation with Expectation-Maximization
Treating $\mathbf{x}^{\rm obs}$ as observed variables while $\mathbf{x}^{\rm mis}$ as latent variables, with the estimated density of the complete data distribution being parameterized as $p_{\bm{\theta}}(\mathbf{x})$, we can formulate the missing value imputation problem using the Expectation-Maximization (EM) algorithm. Specifically, when the complete data distribution $p_{\bm{\theta}}(\mathbf{x})$ is available, the optimal estimation of missing values is given by $\mathbf{x}^{{\rm mis}*}={\mathbb{E}}_{\mathbf{x}^{\rm mis}}p(\mathbf{x}^{\rm mis}|\mathbf{x}^{\rm obs},\bm{\theta})$. Conversely, when the missing entries $\mathbf{x}^{\rm mis}$ are known, the density parameters can be optimized via maximum likelihood estimation: $\bm{\theta}^{*}=\mathop{\arg\max}\limits_{\bm{\theta}}p(\mathbf{x}|\bm{\theta})$. Consequently, with the initial estimation of missing values $\mathbf{x}^{\rm mis}$, the model parameters $\bm{\theta}$ and the missing value $\mathbf{x}^{\rm mis}$ can be optimized by iteratively applying M-step and E-step:
  • Maximization-step: Fix $\mathbf{x}^{\rm mis}$, update $\bm{\theta}^{*}=\mathop{\arg\max}\limits_{\bm{\theta}}p(\mathbf{x}|\bm{\theta})=\mathop{\arg\max}\limits_{\bm{\theta}}p_{\bm{\theta}}(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$.
  • Expectation-step: Fix $\bm{\theta}$, update $\mathbf{x}^{\rm mis*}={\mathbb{E}}_{\mathbf{x}^{\rm mis}}p(\mathbf{x}^{\rm mis}|\mathbf{x}^{\rm obs},\bm{\theta})$.
Remark 1.
The Expectation step can be equivalently rewritten as $\mathbf{x}^{*}={\mathbb{E}}_{\mathbf{x}}\;p(\mathbf{x}|\mathbf{x}^{\rm obs},\bm{\theta})$, where $\mathbf{x}^{*}=(\mathbf{x}^{{\rm obs}},\mathbf{x}^{{\rm mis}*})$. In the remaining sections, we use $\mathbf{x}^{*}$ to denote the E-step for convenience.
4 Methodology
In this section, we introduce DiffPuter- Iterative Missing Data Imputation with Diffusion. Based on the Expectation-Maximization (EM) algorithm, DiffPuter updates the density parameter $\bm{\theta}$ and hidden variables $\mathbf{x}^{\rm mis}$ in an iterative manner. Fig. 1 shows the overall architecture and training process of DiffPuter: 1) The M-step fixes the missing entries $\mathbf{x}^{\rm mis}$, then a diffusion model is trained to estimate the density of the complete data distribution $p_{\bm{\theta}}(\mathbf{x})=p_{\bm{\theta}}(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$; 2) The E-step fixes the model parameters $\bm{\theta}$, then we update the missing entries $\mathbf{x}^{\rm mis}$ via the reverse process of the learned diffusion model $p_{\bm{\theta}}(\mathbf{x})$. The above two steps are executed iteratively until convergence. The following sections introduce the M-step and E-step of DiffPuter, respectively. To avoid confusion, we use $\mathbf{x},\mathbf{x}_{t},\mathbf{x}^{\rm obs}$, etc. to denote samples from real data, $\tilde{\mathbf{x}},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{x}}^{\rm obs}$, etc. to denote samples obtained by the model $\bm{\theta}$, while $\hat{\mathbf{x}},\hat{\mathbf{x}}_{t},\hat{\mathbf{x}}_{\rm obs}$, etc. to denote the specific values of variables.
Figure 1: An overview of the architecture of the proposed DiffPuter. DiffPuter utilizes analog bit encoding to transform discrete variables into continuous ones and use the mean of observed values to initialize the missing entries. The EM algorithm alternates the process of 1) fixing $\mathbf{x}^{\rm mis}$ and estimate diffusion model parameter $\mathbf{\bm{\theta}}$, 2) fixing $\mathbf{\bm{\theta}}$ and estimate $\mathbf{x}^{\rm mis}$, for $K$ iterations. The final imputation result $\mathbf{x}^{*}$ is returned from the E-step of the last iteration.
4.1 M-step: Density Estimation with Diffusion Models
Given an estimation of complete data $\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$, M-step aims to learn the density of $\mathbf{x}$, parameterized by model $\bm{\theta}$, i.e., $p_{\bm{\theta}}(\mathbf{x})$. Inspired by the impressive generative modeling capacity of diffusion models [27, 10], DiffPuter learns $p_{\bm{\theta}}(\mathbf{x})$ through a diffusion process, which consists of a forward process that gradually adds Gaussian noises of increasing scales to $\mathbf{x}$, and a reverse process that recovers the clean data from the noisy one:
[…]
In the forward process, $\mathbf{x}_{0}=\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$ is the currently estimated data at time 00, and $\mathbf{x}_{t}$ is the diffused data at time $t$. $\sigma(t)$ is the noise level, i.e., the standard deviation of Gaussian noise, at time $t$. The forward process has defined a series of data distribution $p(\mathbf{x}_{t})=\int_{\mathbf{x}_{0}}p(\mathbf{x}_{t}|\mathbf{x}_{0})p(\mathbf{x}_{0}){\rm d}\mathbf{x}_{0}$, and $p(\mathbf{x}_{0})=p(\mathbf{x})$. Note that when restricting the mean of $\mathbf{x}_{0}$ as 00 and the variance of $\mathbf{x}_{0}$ be small (e.g., via standardization), $p(\mathbf{x}_{t})$ approaches a tractable prior distribution $\pi(\mathbf{x})$ at $t=T$ when $\sigma(T)$ is large enough, meaning $p(\mathbf{x}_{T})\approx\pi(\mathbf{x})$ [26]. In our formulation in Eq. 1, $p(\mathbf{x}_{T})\approx\pi(\mathbf{x})={\mathcal{N}}(\bm{0},\sigma^{2}(T)\mathbf{I})$.
In the reverse process, $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ is the gradient of $\mathbf{x}_{t}$’s log-probability w.r.t., to $\mathbf{x}_{t}$, and is also known as the score function. $\bm{\omega}_{t}$ is the standard Wiener process. The model is trained by (conditional) score matching [27], which essentially utilizes a neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ to approximate the conditional score-function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})$, which approximates $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ in expectation:
${\mathcal{L}}_{\rm SM}={\mathbb{E}}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0})}{\mathbb{E}}_{t\sim p(t)}{\mathbb{E}}_{\bm{\varepsilon}\sim{\mathcal{N}}(\bm{0},\mathbf{I})}\|\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})\|_{2}^{2},\;~{}\text{where}~{}\mathbf{x}_{t}=\mathbf{x}_{0}+\sigma(t)\bm{\varepsilon}.$ (3)
Since the score of conditional distribution has analytical solutions, i.e., $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})=\frac{\nabla_{\mathbf{x}_{t}}p(\mathbf{x}_{t}|\mathbf{x}_{0})}{p(\mathbf{x}_{t}|\mathbf{x}_{0})}=-\frac{\mathbf{x}_{t}-\mathbf{x}_{0}}{\sigma^{2}(t)}=-\frac{\bm{\varepsilon}}{\sigma(t)}$, Eq. 3 can be interpreted as training a neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ to approximate the scaled noise. Therefore, $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ is also known as the denoising function, and in this paper, it is implemented as a five-layer MLP (see Appendix C.4).
Remark 2.
Starting from the prior distribution $\mathbf{x}_{T}\sim\pi(\mathbf{x})$, and apply the reverse process in Eq. 2 with $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ replaced with $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$, we obtain a series of distributions $p_{\bm{\theta}}(\mathbf{x}_{t}),\forall t\in[0,T]$
Remark 3.
(Corollary 1 in [26]) Let $p(\mathbf{x})=p(\mathbf{x}_{0})$ be the data distribution, and $p_{\bm{\theta}}(\mathbf{x})=p_{\bm{\theta}}(\mathbf{x}_{0})$ be the marginal data distribution obtained from the reverse process, then the score-matching loss function in Eq. 3 is an upper bound of the negative-log-likelihood of real data ${\bm{x}}\sim p({\bm{x}})$ over $\bm{\theta}$. Formally,
$-{\mathbb{E}}_{p(\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x})]\leq{\mathcal{L}}_{\rm SM}(\bm{\theta})+Const,$ (4)
where $Const$ is a constant independent of $\bm{\theta}$.
Remark 3 indicates that the $\bm{\theta}$ optimized by Eq. 3 is the maximum likelihood estimation of data distribution $p(\mathbf{x})$. Consequently, $p_{\bm{\theta}}({\bm{x}})$ approximates accurately $p(\mathbf{x})$ with enough capacity. The algorithmic illustration of DiffPuter’s M-step is summarized in Algorithm 1.
4.2 E-Step: Missing Data Imputation with a Learned Diffusion Model
Given the current estimation of data distribution $p_{\bm{\theta}}(\mathbf{x})$, the E-step aims to obtain the distribution of complete data, conditional on the observed values, i.e., $p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs})$, such that the estimated complete data $\mathbf{x}^{*}$ can be updated by taking the expectation, i.e., $\mathbf{x}^{*}={\mathbb{E}}_{\mathbf{x}}\;p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs})$.
When there is an explicit density function for $p_{\bm{\theta}}(\mathbf{x})$, or when the conditional distribution $p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs})$ is tractable (e.g., can be sampled), computing ${\mathbb{E}}_{\mathbf{x}}\;p(\mathbf{x}|\mathbf{x}^{\rm obs},\bm{\theta})$ becomes feasible. While most deep generative models such as VAEs and GANs support convenient unconditional sampling from $p_{\bm{\theta}}(\mathbf{x})$, they do not naturally support conditional sampling, e.g., $\tilde{\mathbf{x}}\sim p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs})$. Luckily, since the diffusion model preserves the size and location of features in both the forward diffusion process and the reverse denoising process, it offers a convenient and accurate solution to perform conditional sampling $p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs})$ from an unconditional model $p_{\bm{\theta}}(\mathbf{x})=p_{\bm{\theta}}(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$.
Specifically, let ${\mathbf{x}}$ be the data to impute, $\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}$ be the values of observed entries, $\mathbf{m}$ be the $0/1$ indicators of the location of missing entries, and $\tilde{\mathbf{x}}_{t}$ be the imputed data at time $t$. Then, we can obtain the imputed data at time $t-\Delta t$ via combining the observed entries from the forward process on ${\mathbf{x}}_{0}=\mathbf{x}$, and the missing entries from the reverse process on $\tilde{\mathbf{x}}_{t}$ from the prior step $t$ [16, 27]:
[…]
Based on the above process, starting from a random noise from the maximum time $T$, i.e., $\tilde{\mathbf{x}}_{T}\sim{\mathcal{N}}(\bm{0},\sigma^{2}(T)\mathbf{I})$, we are able to obtain a reconstructed $\tilde{\mathbf{x}}_{0}$, such that the observed entries of $\tilde{\mathbf{x}}_{0}$ is the same as those of $\mathbf{x}_{0}$, i.e., $\tilde{\mathbf{x}}^{\rm obs}_{0}=\hat{\mathbf{x}}^{\rm obs}_{0}$. In Theorem 1, we prove that via the algorithm above, the obtained $\tilde{\mathbf{x}}_{0}$ is sampled from the desired conditional distribution, i.e., $\tilde{\mathbf{x}}_{0}\sim p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$.
Theorem 1.
Let $\tilde{\mathbf{x}}_{T}$ be a sample from the prior distribution $\pi(\mathbf{x})={\mathcal{N}}(\mathbf{0},\sigma^{2}(T)\mathbf{I})$, $\mathbf{x}$ be the data to impute, and the known entries of $\mathbf{x}$ are denoted by $\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}$. The score function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ is approximated by neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ Applying Eq. 5, Eq. 6, and Eq. 7 iteratively from $t=T$ until $t=0$, then $\tilde{\mathbf{x}}_{0}$ is a sample from $p_{\bm{\theta}}(\mathbf{x})$, under the condition that its observed entries $\tilde{\mathbf{x}}_{0}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}$. Formally,
$\tilde{\mathbf{x}}_{0}\sim p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$ (8)
See proof in Appendix A.1. Theorem 1 demonstrates that with a learned diffusion model $\bm{\theta}$, we are able to obtain samples exactly from the conditional distribution $p_{\bm{\theta}}(\mathbf{{\bm{x}}}|\mathbf{x}^{\rm obs})$ through the aforementioned imputation process. Therefore, it recovers the E-step in the EM algorithm.
Discretization.
The above imputing process involves recovering $\tilde{\mathbf{x}}_{t}$ continuously from $t=T$ to $t=0$, which is infeasible in practice. In implementation, we discretize the process via $M+1$ discrete descending timesteps $t_{M},t_{M-1},\cdots,t_{i},\cdots,t_{0}$, where $t_{M}=T,t_{0}=0$. Therefore, starting from $\tilde{\mathbf{x}}_{t_{M}}\sim p_{\bm{\theta}}(\mathbf{x}_{t_{M}}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$, we obtain $p_{\bm{\theta}}(\mathbf{x}_{t_{i}}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$ from $i=M-1$ to $i=0$.
Since the desired imputed data $\mathbf{x}^{{\rm mis}*}$ is the expectation, i.e., $\mathbf{x}^{{\rm mis}*}={\mathbb{E}}_{\mathbf{x}^{\rm mis}}p(\mathbf{x}^{\rm mis}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$. We sample $\tilde{\mathbf{x}}_{0}^{\rm mis}$ for $N$ times, and take the average value as the imputed $\mathbf{x}^{{\rm mis}*}$. The algorithmic illustration of DiffPuter’s E-step is summarized in Algorithm 2.
Input: Data $\mathbf{x}\in p(\mathbf{x})\in{\mathbb{R}}^{1\times d}$.
1
Output: Score network $\bm{\epsilon}_{\bm{\theta}}$
2
3# $\bm{\varepsilon}\sim{\mathcal{N}}(\bm{0},\mathbf{I})\in{\mathbb{R}}^{d}$
4 while not converging do
5       $\mathbf{x}_{0}=\mathbf{x}$
6       $t\sim p(t)$
7       $\mathbf{x}_{t}=\mathbf{x}_{0}+\sigma(t)\cdot\bm{\varepsilon}$
8       $\ell(\bm{\theta})=\|\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)-\frac{-\bm{\varepsilon}}{\sigma(t)}\|_{2}^{2}$
9       Update $\bm{\theta}$ via Adam optimizer
10
Algorithm 1 M-step: Density Estimation using Diffusion Model
Input: Score network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$. Data with missing values ${\mathbf{x}}\in{\mathbb{R}}^{d}$, mask $\mathbf{m}\in\{0,1\}^{d}$
1
Output: Imputed data $\mathbf{x}^{*}$
2
3Sample $\tilde{\mathbf{x}}_{t_{M}}\sim{\mathcal{N}}(\bm{0},\sigma^{2}(t_{M})\mathbf{I})$
4 for $j\leftarrow$ $1$ to $N$ do
5       for $i\leftarrow$ $M$ to $1$ do
6             $\mathbf{x}_{t_{i-1}}^{\rm forward}=\mathbf{x}+\sigma(t_{i-1})\cdot\bm{\varepsilon}$
7             $\mathbf{x}_{t_{i-1}}^{\rm reverse}=\tilde{\mathbf{x}}_{t_{i}}+\int_{t_{i}}^{t_{i-1}}{\rm d}\tilde{\mathbf{x}}_{t_{i}}$
8             $\tilde{\mathbf{x}}_{t_{i-1}}=\mathbf{m}\odot\mathbf{x}_{t_{i-1}}^{\rm forward}+(1-\mathbf{m})\odot\mathbf{x}_{t_{i-1}}^{\rm reverse}$
9
10
[…]
Algorithm 2 E-step: Missing Data Imputation
4.3 Practice of DiffPuter
Input: Data $\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})\sim p(\mathbf{x})$. Number of iterations $K$.
1
Output: Imputed data at $K$-th iteration $\mathbf{x}^{(K)}$
2 # Analog bit encoding
3 $\mathbf{x}\leftarrow{\rm AnaBit}(\mathbf{x})$
4 $\mathbf{x}^{{\rm obs}(0)}\leftarrow{\rm Standardize}(\mathbf{x}^{{\rm obs}})$
5 $\mathbf{x}^{{\rm mis}(0)}=\bm{0}$
6 $\mathbf{x}^{(0)}=(\mathbf{x}^{{\rm obs}(0)},\mathbf{x}^{{\rm mis}(0)})$
7 for $k\leftarrow$ $1$ to $K$ do
8       $\bm{\theta}^{(k)}=$ M-step $(\mathbf{x}^{(k-1)})$
9       $\mathbf{x}^{(k)}=$ E-step $(\mathbf{x}^{(k-1)},\bm{\theta}^{(k)})$
10
11Return $\mathbf{x}^{(K)}=(\mathbf{x}^{{\rm obs}(K)},\mathbf{x}^{{\rm mis}(K)})$
Algorithm 3 DiffPuter for Missing Data Imputation.
Analog bits for Discrete Variables.
A tabular dataset typically encompasses both numerical (i.e., continuous) data and categorical (i.e., discrete) data. However, conventional diffusion models assume continuous Gaussian noise, thereby rendering them inapplicable directly to categorical variables. To this end, for a column of discrete variables, we encode each category into its corresponding binary code based on its index. A discrete variable with $C$ possible values is encoded into a $\lceil\log_{2}C\rceil$-dimensional vector. For example, when $C=14$, and we obtain a discrete value with index $5=2^{2}+2^{0}$, we transform it into a $4$-dimensional vector $[0,1,0,1]$ , which are treated as float variables in computation. When recovering the categorical variable from the imputed encoding, we set the corresponding entries in the vector to $1$ if they exceed $0.5$, otherwise 00. Finally, we convert the binary 0/1 vector to its corresponding categorical value.
Preprocessing and Initialization.
Since different columns of the tabular data have distinct meanings and might have distinct scales, following existing literature, we compute the mean and standard deviation of each column’s observed values. Then, each column is standardized with zero mean and unit standard deviation. In addition, the execution of the EM algorithm requires the initialized values of missing entries $\mathbf{x}^{\rm mis(0)}$, which might have a huge impact on the model’s convergence. For simplicity, we initialize the missing entries of each column to be the mean of the column’s observed values, equivalent to setting $\mathbf{x}^{\rm mis(0)}=0$ everywhere (since the data has been standardized).
Iterative Training.
To obtain a more accurate estimation of complete data $\mathbf{x}$, DiffPuter iteratively executes the M-step and E-step. To be specific, let $\mathbf{x}^{(k)}=(\mathbf{x}^{{\rm obs}(k)},\mathbf{x}^{{\rm obs}(k)})=(\mathbf{x}^{{\rm obs}},\mathbf{x}^{{\rm obs}(k)})$ be the estimation of complete data at $k$-th iteration, $\bm{\theta}^{(k)}$ be the diffusion model’s parameters at $k$-th iteration, we write $\bm{\theta}^{(k+1)}$ as a function of $\mathbf{x}^{(k)}$, i.e., $\bm{\theta}^{(k+1)}=$ M-step $(\mathbf{x}^{(k)})$, and $\mathbf{x}^{(k+1)}$ as a function of $\bm{\theta}^{(k+1)}$ and $\mathbf{x}^{(k)}$, i.e., $\mathbf{x}^{(k+1)}=$ E-step $(\mathbf{x}^{(k)},\bm{\theta}^{(k+1)})$. Therefore, with the initialized $\mathbf{x}^{(0)}$ and the maximum iteration $K$, we are able to obtain $\mathbf{x}^{(k)}$ from $k=1$ to $K$.
Combining the above designs, we present the overall description of DiffPuter in Algorithm 3.
5 Experiments
In this section, we conduct experiments to study the efficacy of the proposed DiffPuter in missing data imputation tasks. The code is available at: https://github.com/hengruizhang98/DiffPuter.
5.1 Experimental Settings
Datasets.
We evaluate the proposed DiffPuter on ten real-world datasets of varying scales that are publicly available. We consider five datasets of only continuous features: California, Letter, Gesture, Magic, and Bean, and five datasets of both continuous and discrete features: Adult, Default, Shoppers, Beijing, and News. The detailed information of these datasets is presented in Appendix C.2. Following previous works [19, 33], we study three missing mechanisms: 1) For missing completely at random (MCAR), the mask for each dimension of every sample is independently sampled from a Bernoulli distribution. 2) For missing at random (MAR), we first select the columns that have no missing values and then employ a logistic model that takes these non-missing columns as input to generate the missing entries in the remaining columns. 3) For MNAR, the masks are generated by a logistic model taking the masked inputs from MCAR. Differences between the three settings are in Appendix C.3. In the main experiments, we set the missing rate as $r=30\%$. For each dataset, we generate $10$ masks according to the missing mechanism and report the mean and standard deviation of the imputing performance. In this section, we only report the performance in the MCAR setting, while the results of the other two settings are in Appendix D.2.
Baselines.
We compare DiffPuter with $16$ powerful imputation methods from different categories: 1) Distribution-matching methods based on optimal transport, including TDM [33] and MOT [19]. 2) Graph-based imputation methods, including GRAPE [31]: a pure bipartite graph-based framework for data imputation, and IGRM [35]: a graph-based imputation method that iteratively reconstructs the friendship network. 3) Iterative methods, including EM with multivariate Gaussian priors [5], MICE [29], MIRACLE [14], SoftImpute [7], and MissForest [28]. 4) Deep generative models, including MIWAE [18], GAIN [30], MCFlow [25], MissDiff [21] and TabCSDI [34]. It is worth noting that MissDiff and TabCSDI are also based on diffusion. We also compare with two recent SOTA imputing methods ReMasker [3] and HyperImpute [8].
Evaluation Protocols.
For each dataset, we use $70\%$ as the training set, and the remaining $30\%$ as the testing set. All methods are trained on the training set. The imputation is applied to both the missing values in the training set and the testing set. Consequently, the imputation of the training set is the ’in-sample’ setting, while imputing the testing set is the ’out-of-sample’ setting. The performance of DiffPuter is evaluated by the divergence between the predicted values and ground-truth values of missing entries. For continuous features, we use Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), while for discrete features, we use Accuracy. The implementation details and hyperparameter settings are presented in Appendix C.4.
Figure 2: MCAR: MAE and RMSE in in-sample imputation tasks. A blank column indicates that the method fails or gets out-of-memory for that dataset. DiffPuter outperforms the most competitive baseline method by $8.10\%$ (MAE score) and $5.64\%$ (RMSE score) by average. Comparison with the entire set of baselines is shown in Fig. 7 in Appendix D.1.
5.2 Main Results
Performance in In-Sample Imputation.
We first evaluate DiffPuter’s performance in the in-sample imputation task. Fig. 2 compares the performance of different imputation methods regarding MAE and RMSE. Due to the large number of baselines, we only present $9$ recent methods out of the $16$ baselines in Fig. 2, and we present the full comparison in Fig. 7 in Appendix D.1. We have the following observations: 1) Across all datasets, DiffPuter provides high-quality imputation results, matching the best methods on some datasets and significantly outperforming the second-best methods on the remaining datasets. 2) The other two diffusion-based methods, MissDiff and TabCSDI, fail to give satisfying imputation performance. 3) The basic EM algorithm with Multivariate Gaussian assumptions demonstrates competitive performance on several datasets, indicating that for relatively simple numerical datasets, basic ML algorithms are not inferior to deep learning methods. 4) Compared to predictive imputation methods, DiffPuter exhibits larger standard deviations because predictive imputation is deterministic, whereas DiffPuter requires stochastic sampling.
Fig. 3 further compares the performance regarding Accuracy for discrete columns. Compared to continuous features, DiffPuter performs on par with SOTA methods for categorical features, not demonstrating a significant advantage as it does with continuous features. Overall, for datasets with mixed-type features, DiffPuter performs on par with state-of-the-art methods for their discrete columns while achieving significantly better results for their continuous columns. Therefore, DiffPuter is better at handling the task of missing data imputation for mixed-type datasets.
Figure 3: MCAR: Acc. in in-sample imputation.
Performance in Out-of-Sample Imputation.
Next, we compare the results on the out-of-sample imputation tasks. Considering that some methods cannot be used for out-of-sample imputation, we reduce the number of compared methods. Additionally, for MOT [19], we employ the Round-Robin method proposed in its paper to adapt it for out-of-sample imputation tasks. Fig. 4 compares the MAEs and RMSEs in the out-of-sample imputation task. Comparing it with the results of in-sample imputation, we can easily observe that some methods exhibit significant performance differences between the two settings. For example, graph-based methods GRAPE and IGRM perform well in in-sample imputation, but their performance degrades significantly in out-of-sample imputation. IGRM even fails on all datasets in the out-of-sample imputation setting. In contrast, our DiffPuter demonstrates similarly excellent performance in both in-sample and out-of-sample imputation. This highlights DiffPuter’s superior performance and robust generalization capabilities.
Figure 4: MCAR: MAE and RMSE in out-of-sample imputation. RMSE on the News dataset are all missing since all imputation methods give rise to abnormal RMSE scores due to extreme values.
5.3 Ablation Studies
In this section, we conduct ablation studies to evaluate the efficacy of the individual designs of DiffPuter, e.g., iterative training, as well as the robustness of DiffPuter, e.g., performance under varying missing ratios.
Does iterative training improve Diffusion Imputation?
In Fig. 6, we present the performance of DiffPuter’s imputation results from increasing EM iterations. Note that $k=0$ represents the imputation result of a randomly initialized denoising network, and $k=1$ represents the performance of a pure diffusion model without iterative refinement. It is clearly observed that a single diffusion imputation achieves only suboptimal performance, while DiffPuter’s performance steadily improves as the number of iterations increases. Additionally, we observe that DiffPuter does not require a large number of iterations to converge. In fact, $4$ to $5$ iterations are sufficient for DiffPuter to converge to a stable and satisfying state.
Figure 5: Performance at different iterations.
Figure 6: Impacts of missing ratios.
Performance under varying missing ratios.
Finally, we investigate if the performance of DiffPuter will be highly impacted by the missing ratios compared with other imputation methods. In Fig. 6, we compare DiffPuter’s MAE score on the Beijing dataset’s continuous features with other SOTA imputation methods under missing ratio $r=0.3,0.5,$ and $0.7$. We find that the most competitive baseline method, Remasker, performs well under low missing ratios. However, when the missing ratio becomes large, its performance drops significantly, being surpassed by all other methods. Our DiffPuter exhibits the most robust performance across increasing missing ratios. Even with a missing ratio of $0.7$, it maintains very good imputation performance.
6 Conclusions
In this paper, we have proposed DiffPuter for missing data imputation. DiffPuter is an iterative method that combines the Expectation-Maximization algorithm and diffusion models, where the diffusion model serves as both the density estimator and missing data imputer. We demonstrate theoretically that the training and sampling process of a diffusion model precisely corresponds to the M-step and E-step of the EM algorithm. Therefore, we can iteratively update the density of the complete data and the values of the missing data. Extensive experiments have demonstrated the efficacy of the proposed method.
Appendix A Proofs
A.1 Proof for Theorem 1
Proof.
First, it is obvious that the observed entries from Eq. 7, i.e., $\tilde{\mathbf{x}}_{t}^{\rm obs}$, satisfy
$\tilde{\mathbf{x}}_{t}^{\rm obs}\sim p({\mathbf{x}}_{t}^{\rm obs}|{\mathbf{x}}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}),\;\forall t.$ (9)
Then, we introduce the following Lemma.
Lemma 1.
Let $\tilde{\mathbf{x}}_{t}\sim p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$. If $\tilde{\mathbf{x}}_{t-\Delta t}$ is a sample obtained from Eq. 5, Eq. 6, and Eq. 7, then $\tilde{\mathbf{x}}_{t-\Delta t}\sim p_{\bm{\theta}}(\mathbf{x}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$, when $\Delta t\rightarrow 0^{+}$.
Proof.
First, note that $\mathbf{x}_{t}$ ($\mathbf{x}_{t-\Delta t}$) is obtained via adding dimensionally-independent Gaussian noises to $\mathbf{x}_{0}=\mathbf{x}$ (see the forward process in Eq. 1), we have
[…]
Therefore, $\tilde{\mathbf{x}}_{t-\Delta t}$’s observed entries $\tilde{\mathbf{x}}_{t-\Delta t}^{\rm obs}$ is sampled from distribution $p(\mathbf{x}^{\rm obs}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$. Then, we turn to the missing entries, where we have the following derivation:
$\begin{split}p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})&=\int p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}_{t},\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}){\rm d}\mathbf{x}_{t}\\ &={\mathbb{E}}_{p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})}p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}_{t},\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})\\ &\approx{\mathbb{E}}_{p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})}p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}_{t})\\ &\approx p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\tilde{\mathbf{x}}_{t}),\\ \end{split}$ (11)
where $\tilde{\mathbf{x}}_{t}$ is a random sample from $p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}^{\rm obs}_{0})$. The first $\approx$’ holds because when $\Delta t\rightarrow 0$, $\mathbf{x}^{\rm mis}_{t-\Delta t}$ is almost predictable via $\mathbf{x}_{t}$ without $\mathbf{x}_{0}^{\rm obs}$. The second $\approx$’ is due to Monte Carlo estimation. Therefore, when $\tilde{\mathbf{x}}_{t}\sim p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$, $p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$ is approximately tractable via $p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\tilde{\mathbf{x}}_{t})$.
Note that $\mathbf{x}_{t-\Delta t}^{\rm mis}$ denotes the missing entries of $\mathbf{x}_{t-\Delta t}$. According to the reverse denoising process in Eq. 2, given $\mathbf{x}_{t}=\tilde{\mathbf{x}}_{t}$, $\mathbf{x}_{t-\Delta t}$ is obtained via integrating ${\rm d}\mathbf{x}_{t}$ from $t$ to $t-{\Delta t}$, i.e.,
$\mathbf{x}_{t-\Delta t}=\tilde{\mathbf{x}}_{t}+\int_{t}^{t-\Delta t}{\rm d}\mathbf{x}_{t}.$ (12)
Therefore, $\tilde{\mathbf{x}}^{\rm mis}_{t-\Delta t}={\mathbf{x}}^{\rm mis}_{t-\Delta t}$ obtained from Eq. 6 is a sample from $p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\tilde{\mathbf{x}}_{t})$. And, approximately, $\tilde{\mathbf{x}}^{\rm mis}_{t-\Delta t}\sim p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$.
Since $\tilde{\mathbf{x}}^{\rm obs}_{t-\Delta t}\sim p(\mathbf{x}^{\rm obs}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}),\tilde{\mathbf{x}}^{\rm mis}_{t-\Delta t}\sim p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$, we have $\tilde{\mathbf{x}}_{t-\Delta t}\sim p_{\bm{\theta}}(\mathbf{x}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$, according to Eq. 10. Therefore, the proof for Lemma 1 is completed. ∎
With Lemma 1, we are able to prove Theorem 1 via induction, as long as $\tilde{\mathbf{x}}_{T}$ is also sampled from $p(\mathbf{x}_{T}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$. This holds because $p(\mathbf{x}_{T}|\mathbf{x}_{0})\approx p(\mathbf{x}_{T})$, given the condition that $\mathbf{x}_{0}=\mathbf{x}$ has zero mean and unit variance, and $\sigma(T)\gg 1$.
Note that the score function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ in Eq. 2 is intractable, it is replaced with the output of the score neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$. Therefore, the finally obtained distribution can be rewritten as $p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$. Therefore, the proof for Theorem  1 is complete.
Appendix B Further explanation of the Diffusion Model
The diffusion model we adopt in Section 4.1 is actually a simplified version of the Variance-Exploding SDE proposed in  [27].
Note that  [27] has provided a unified formulation via the Stochastic Differential Equation (SDE) and defines the forward process of Diffusion as
${\rm d}\mathbf{x}={\bm{f}}(\mathbf{x},t){\rm d}t+g(t)\;{\rm d}{\bm{w}}_{t},$ (13)
where ${\bm{f}}(\cdot)$ and $g(\cdot)$ are the drift and diffusion coefficients and are selected differently for different diffusion processes, e.g., the variance preserving (VP) and variance exploding (VE) formulations. $\bm{\omega}_{t}$ is the standard Wiener process. Usually, $f(\cdot)$ is of the form ${\bm{f}}(\mathbf{x},t)=f(t)\;\mathbf{x}$. Thus, the SDE can be equivalently written as
${\rm d}\mathbf{x}=f(t)\;\mathbf{x}\;{\rm d}t+g(t)\;{\rm d}{\bm{w}}_{t}.$ (14)
Let $\mathbf{x}$ be a function of the time $t$, i.e., $\mathbf{x}_{t}=\mathbf{x}(t)$, then the conditional distribution of $\mathbf{x}_{t}$ given $\mathbf{x}_{0}$ (named as the perturbation kernel of the SDE) could be formulated as:
$p(\mathbf{x}_{t}|\mathbf{x}_{0})={\mathcal{N}}(\mathbf{x}_{t};s(t)\mathbf{x}_{0},s^{2}(t)\sigma^{2}(t){\bm{I}}),$ (15)
where
$s(t)=\exp\left(\int_{0}^{t}f(\xi){\rm d}\xi\right),\;\mbox{and}\;\sigma(t)=\sqrt{\int_{0}^{t}\frac{g^{2}(\xi)}{s^{2}(\xi)}{\rm d}\xi}.$ (16)
Therefore, the forward diffusion process could be equivalently formulated by defining the perturbation kernels (via defining appropriate $s(t)$ and $\sigma(t)$).
Variance Exploding (VE) implements the perturbation kernel Eq. 15 by setting $s(t)=1$, indicating that the noise is directly added to the data rather than weighted mixing. Therefore, The noise variance (the noise level) is totally decided by $\sigma(t)$. The diffusion model used in DiffPuter belongs to VE-SDE, but we use linear noise level (i.e., $\sigma(t)=t$) rather than $\sigma(t)=\sqrt{t}$ in the vanilla VE-SDE [27]. When $s(t)=1$, the perturbation kernels become:
$p(\mathbf{x}_{t}|\mathbf{x}_{0})={\mathcal{N}}(\mathbf{x}_{t};\bm{0},\sigma^{2}(t)\mathbf{I})\;\Rightarrow\;\mathbf{x}_{t}=\mathbf{x}_{0}+\sigma(t)\bm{\varepsilon},$ (17)
which aligns with the forward diffusion process in Eq. 1.
The sampling process of diffusion SDE is given by:
${\rm d}\mathbf{x}=[{\bm{f}}(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})]{\rm d}t+g(t){\rm d}{\bm{w}}_{t}.$ (18)
For VE-SDE, $s(t)=1\Leftrightarrow{\bm{f}}(\mathbf{x},t)=f(t)\cdot\mathbf{x}=\bm{0}$, and
$\begin{split}&\sigma(t)=\sqrt{\int_{0}^{t}g^{2}(\xi){\rm d}\xi}\Rightarrow\int_{0}^{t}g^{2}(\xi){\rm d}\xi=\sigma^{2}(t),\\ &g^{2}(t)=\frac{{\rm d}\sigma^{2}(t)}{{\rm d}t}=2\sigma(t)\dot{\sigma}(t),\\ &g(t)=\sqrt{2\sigma(t)\dot{\sigma}(t)}.\end{split}$ (19)
Plugging $g(t)$ into Eq. 18, the reverse process in Eq. 2 is recovered:
${\rm d}\mathbf{x}_{t}=-2\sigma(t)\dot{\sigma}(t)\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}){\rm d}t+\sqrt{2\sigma(t)\dot{\sigma}(t)}{\rm d}\bm{\omega}_{t}.$ (20)
Appendix C Experimental Details
C.1 Configurations.
We conduct all experiments with:
  • Operating System: Ubuntu 22.04.3 LTS
  • CPU: Intel 13th Gen Intel(R) Core(TM) i9-13900K
  • GPU: NVIDIA GeForce RTX 4090 with 24 GB of Memory
  • Software: CUDA 12.2, Python 3.9.16, PyTorch [22] 1.12.1
C.2 Datasets
Table 1: Statistics of datasets. # Num stands for the number of numerical columns, and # Cat stands for the number of categorical columns.
Dataset # Rows # Num # Cat # Train (In-sample) # Test (Out-of-Sample)
California Housing $20,640$ $9$ $14,303$ $6,337$
Letter Recognition $20,000$ $16$ $14,000$ $6,000$
Gesture Phase Segmentation $9,522$ $32$ $6,665$ $2,857$
Magic Gamma Telescope $19,020$ $10$ $13,314$ $5,706$
Dry Bean $13,610$ $17$ $9,527$ $4,083$
Adult Income $32,561$ $6$ $8$ $22,792$ $9,769$
Default of Credit Card Clients $30,000$ $14$ $10$ $21,000$ $9,000$
Online Shoppers Purchase $12,330$ $10$ $7$ $8,631$ $3,699$
Beijing PM2.5 data $43,824$ $6$ $5$ $29,229$ $12,528$
Online News Popularity $39,644$ $45$ $2$ $27,790$ $11,894$
C.3 Missing mechanisms
According to how the masks $\mathbf{m}$ are generated, there are three mainstream mechanisms of missingness, namely missing patterns: 1) Missing completely at random (MCAR) refers to the case that the probability of an entry being missing is independent of the data, i.e., $p(\mathbf{m}|\mathbf{x})=p(\mathbf{m})$. 2) In missing at random (MAR), the probability of missingness depends only on the observed values, i.e., $p(\mathbf{m}|\mathbf{x})=p(\mathbf{m}|\mathbf{x}^{\rm obs})$ 3) All other cases are classified as missing not at random (MNAR), where the probability of missingness might also depend on other missing entries.
The code for generating masks according to the three missing mechanisms is also provided in the supplementary.
C.4 Implementations and hyperparameters
We use a fixed set of hyperparameters, which will save significant efforts in hyperparameter-tuning when applying DiffPuter to more datasets. For the diffusion model, we set the maximum time $T=80$, the noise level $\sigma(t)=t$, which is linear to $t$. The score/denoising neural network $\bm{\epsilon}(\mathbf{x}_{t},t)$ is implemented as a $5$-layer MLP with hidden dimension $1024$. $t$ is transformed to sinusoidal timestep embeddings and then added to $\mathbf{x}_{t}$, which is subsequently passed to the denoising function. When using the learned diffusion model for imputation, we set the number of discrete steps $M=50$ and the number of sampling times per data sample $N=10$. DiffPuter is implemented with Pytorch, and optimized using Adam [11] optimizer with a learning rate of $1\times 10^{-4}$.
Architecture of denoising neural network.
We use the same architecture of denoising neural network as in two recent tabular diffusion models for tabular data synthesis [13, 32]. The denoising MLP takes the current time step $t$ and the feature vector $\mathbf{x}_{t}\in{\mathbb{R}}^{1\times d}$ as input. First, $\mathbf{x}_{t}$ is fed into a linear projection layer that converts the vector dimension to be $d_{\rm hidden}=1024$:
$\mathbf{h}_{0}=\texttt{FC}_{\rm in}(\mathbf{x}_{t})\in{\mathbb{R}}^{1\times d_{\rm hidden}},$ (21)
where $\mathbf{h}_{0}$ is the transformed vector, and $d_{\rm hidden}$ is the output dimension of the input layer.
Then, following the practice in TabDDPM [13], the sinusoidal timestep embeddings $\mathbf{t}_{\rm emb}\in{\mathbb{R}}^{1\times d_{\rm hidden}}$ is added to $\mathbf{h}_{0}$ to obtain the input vector $\mathbf{h}_{\rm hidden}$:
$\mathbf{h}_{\rm in}=\mathbf{h}_{0}+\mathbf{t}_{\rm emb}.$ (22)
The hidden layers are three fully connected layers of the size $d_{\rm hidden}-2*d_{\rm hidden}-2*d_{\rm hidden}-d_{\rm hidden}$, with SiLU activation functions:
$\begin{split}\mathbf{h}_{1}&=\texttt{SiLU}(\texttt{FC}_{1}(\mathbf{h}_{0})\in{\mathbb{R}}^{1\times 2*d_{\rm hidden}}),\\ \mathbf{h}_{2}&=\texttt{SiLU}(\texttt{FC}_{2}(\mathbf{h}_{1})\in{\mathbb{R}}^{1\times 2*d_{\rm hidden}}),\\ \mathbf{h}_{3}&=\texttt{SiLU}(\texttt{FC}_{3}(\mathbf{h}_{2})\in{\mathbb{R}}^{1\times d_{\rm hidden}}).\\ \end{split}$ (23)
The estimated score is obtained via the last linear layer:
$\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)=\mathbf{h}_{\rm out}=\texttt{FC}_{\rm out}(\mathbf{h}_{3})\in{\mathbb{R}}^{1\times d}.$ (24)
Finally, $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is applied to Eq. 3 for model training.
C.5 Implementations of baselines
We implement most the baseline methods according to the publicly available codebases:
MissDiff [21] does not provide its official implementations. Therefore, we obtain its results based on our own implementation.
The codes for all the methods are available in the supplementary.
Appendix D Additional Results
D.1 Comparison with all baselines in MCAR
In Fig. 7 and Fig. 8, we compare the proposed DiffPuter with all the covered baseline methods under the MCAR setting.
Figure 7: MCAR (full): MAE and RMSE in in-sample imputation tasks. The missing columns indicate that the method fails or gets out-of-memory for that dataset.
Figure 8: MCAR (full): Acc. in in-sample imputation.
D.2 Missing at Random (MAR) and Missing not at Random (MNAR)
In Fig. 9 and Fig. 10, we present the in-sample imputation MAE and RMSE scores of different methods in MAR and MNAR settings, respectively. In general, We observe similar trends as shown in the MCAR setting in Fig. 2, and the performance of the proposed DiffPuter isn’t impacted by different missing mechanisms seriously.
Figure 9: MAR: MAE and RMSE in in-sample imputation tasks. The missing columns indicate that the method fails or gets out-of-memory for that dataset.
Figure 10: MNAR: MAE and RMSE in in-sample imputation tasks. The missing columns indicate that the method fails or gets out-of-memory for that dataset.